3/5z + 7= 11 (I can't remember how to do this!)
Answer by checkley77(12844)

3Z/5+7=11
3Z/5=11-7
3Z/5=4
3Z=4*5
3Z=20
Z=20/3 ANS.
PROOF:
3/5*20/3+7=11
60/15=11-7
4=4